radomes.openskiesproject.org
thehelperforyou
thehelperforyou thehelperforyou
  • 17-01-2017
  • Mathematics
Answered

if the probability of a train being on time is 0.24 and the probability of it being late is 0.66 what is the probability of it being early

Verified Answer :

simplycheza
simplycheza simplycheza
  • 17-01-2017

0.24 + 0.66 = 0.90
1 - 0.90 = 0.10

Answer Link

Related Questions

Binge Eating Disorder is diagnosed by experiencing bingeeating episodes without purging more than times per week for a One two months b Two three months c Two s
what is 57 times 89
a cheetah can run up to 75 miles per hour how many yards per minute is this
what is the answer for question 9
Which of the following leaders was NOT a participant in the Six Day War a President Gamal Nasser of Egypt b King Hussein of Jordan c Saddam Hussein of Iraq d Pr
Name and define 3 of the organelles that we are learning about
What is 216 repeat rewrite as a fraction
Genes areO responsible for inherited traitsOunits of coded DNAlocated in a specific place on a chromosomeO all of the above
1 Lincolns focus in the Gettysburg Address was to 1 point A campaign for a reelection Burge the south to surrender Cfree African American Slaves in the South D
It was a children versus grownups competition at school One event required the adult to throw a basketball as far as he could When it was the childs turn he had